#1.Calculate the [H3O+] of aqueous solution with the pOH = 4.20.
Express your answer using two significant figures.

[H3O+] = ???M

#2. Calculate the [OH−] of aqueous solution with the pOH = 4.20.
Express your answer using two significant figures.

[OH−] = ??? M

#3. Calculate the pH of aqueous solution with the pOH = 4.20.
Express your answer using two decimal places.

pH = ??

State whether the solution is acidic, basic, or neutral:
Acidic solution
Neutral solution
Basic solution
1
Expert's answer
2015-11-20T10:18:32-0500
1. pH=-lg[H3O+]
pOH=-lg[OH-]
Kw=[H3O+][OH-]=10^-14
[H3O+]+[OH-]=14
pH=14-4.20=9.80
[H3O+]=10^-pH
[H3O+]=10^-9.80=1.58·10^-10 M

2. pOH=-lg[OH-]
[OH-]=10^-pOH
[OH-]=10^-4.20=6.31·10^-5 M
The solution is alkaline because pH>7

3. pH=-lg[H3O+]
Kw=[H3O+][OH-]=10^-14
[H3O+]+[OH-]=14
pH=14-4.20=9.80
The solution is alkaline because pH>7